# Dynamic Target Attack

### v2

@@ 5 Conclusion
In this paper, we propose Dynamic Target Attack (DTA), a novel jailbreak method that relies on directly sampling candidate responses as targets to optimize the adversarial prompt. In each optimization round, DTA iteratively samples multiple responses and selects the most harmful one as the current round target. In contrast to existing methods, DTA selects targets from the target model’s relatively high-density region, significantly reducing the discrepancy between the target and the model’s output distribution, thereby accelerating convergence to an effective adversarial prompt. Throughout our extensive experiments on diverse jailbreak benchmarks, DTA demonstrates its effectiveness against multiple recent safety-aligned LLMs under both white- and black-box settings, consistently outperforming the state-of-the-art baselines.
Appendix A Related Work
Here, we introduce six white-box and three black-box methods as our baselines in this paper:
A.1 White-box baselines
GCG [6] (Greedy Coordinate Gradient) is a foundational white-box attack. It employs a greedy, coordinate descent-based optimization to find an adversarial suffix $S$ that, when appended to a harmful prompt $P$, maximizes the probability of the model generating a fixed, predefined target response (e.g., “Sure, here is…”).
COLD-Attack. To address the limitations of high-perplexity attacks like GCG, Guo et al. (2024) introduced COLD-Attack, a framework that formally reframes controllable jailbreaking as a controllable text generation problem. By adapting the COLD algorithm, this method operates in the continuous logit space rather than on discrete tokens. It defines a composite energy function to unify multiple differentiable constraints, such as attack success, fluency, and semantic similarity. The framework uses Langevin dynamics to sample optimal logits that balance these competing objectives, and then employs a specialized LLM-guided decoding process to ensure the final text is coherent. This unified approach enables the generation of highly stealthy, low-perplexity attacks and introduces novel, controllable scenarios like paraphrasing and position-constrained attacks, which are not achievable by previous methods.
llm-adaptive [10] is an adaptive methodology for efficiently generating successful jailbreaking attacks against state-of-the-art, safety-aligned large language models. Operating under gray-box or black-box assumptions, its core strategy combines customized, handcrafted prompt templates with simple optimization algorithms, like random search, to iteratively refine attack strings based on model feedback. A key component of its effectiveness is the "self-transfer" technique, where successful attack patterns discovered on one harmful behavior are used to rapidly accelerate the attack on others, and it further exploits model-specific vulnerabilities, such as optimizing against token log-probabilities or leveraging unique API features like prefilling, to bypass safety filters with surprisingly few queries.
I-GCG. Jia et al. [13] proposed I-GCG, enhancing optimization-based jailbreaks by addressing the limitations of GCG. They found that GCG’s simple target template (e.g., "Sure, here is…") was insufficient and instead introduced diverse target templates containing harmful self-suggestions (e.g., "my output is harmful") to more effectively mislead the LLM. To improve optimization efficiency, they developed an automatic multi-coordinate updating strategy that adaptively replaces multiple tokens per step, accelerating convergence. Additionally, they employed an easy-to-hard initialization technique, using a suffix generated from a simple malicious query as the starting point for more difficult ones.
AdvPrefix [8] is an optimization objective used in gradient-based adversarial attacks like GCG to "jailbreak" Large Language Models (LLMs). The goal is to find a short, often nonsensical, string of characters (the "prefix") that, when attached to a user’s prompt, hijacks the model’s output generation. This objective function is optimized to find the prefix that maximizes the probability of the LLM generating a specific, malicious target response (e.g., harmful instructions or hate speech), thereby bypassing its safety alignments.
RLbreaker [9] is an automated jailbreak prompt generation method that leverages Deep Reinforcement Learning (DRL) to systematically bypass the safety alignments of Large Language Models (LLMs). The system features three components: a DRL agent (attacker) that iteratively generates and refines prompts, a target LLM (victim), and an evaluator LLM (judge) that provides a reward signal based on the attack’s success. During training, the agent continuously modifies a prompt (up to 5 steps) and attacks the target, learning from the evaluator’s feedback through this trial-and-error process to master strategies for creating effective jailbreak prompts.
A.2 Black-box baselines
ReNeLLM. To address the limitations of manual and optimization-based jailbreaks, Ding et al. [18] proposed ReNeLLM, an automated framework that leverages LLMs themselves to generate effective and transferable jailbreak prompts. Their method generalizes the attack into two main stages: (1) Prompt Rewriting, which automatically applies a series of linguistic transformations (such as paraphrasing, misspelling sensitive words, or partial translation) to disguise the harmful intent of an initial query, and (2) Scenario Nesting, which embeds the rewritten prompt into a seemingly benign and common instruction-following scenario, such as code completion, table filling, or text continuation. This combined approach is shown to effectively shift the target LLM’s execution priority from safety to usefulness, enabling it to bypass safeguards and achieve high attack success rates on state-of-the-art models like GPT-4 and Claude-2.
PAP. Zeng et al. [19] introduce a novel jailbreak methodology, Persuasive Adversarial Prompts (PAPs), which reframes the attack vector by treating LLMs as human-like communicators susceptible to social influence rather than as mere instruction followers. Their approach is systematically grounded in a comprehensive taxonomy of 40 persuasion techniques derived from decades of social science research. The method employs a two-stage framework: first, they train a "Persuasive Paraphraser" by fine-tuning an LLM (e.g., GPT-3.5); second, this trained model is deployed to automatically generate new, human-readable PAPs at scale. The effectiveness of these prompts in bypassing safety alignments is then systematically evaluated across various LLMs, including Llama 2 and the GPT series, using a GPT-4 Judge to classify the harmfulness of the outputs.
TAP. Mehrotra et al. [20] propose Tree of Attacks with Pruning (TAP), an automated, black-box method for generating interpretable jailbreaks for LLMs. TAP utilizes an attacker LLM to iteratively generate and refine candidate prompts in a tree structure (branching) and an evaluator LLM to assess them. This process involves two pruning phases: first, pruning off-topic prompts before querying the target, and second, pruning low-scoring prompts after assessing the target’s response.
Appendix B Supplementary Materials on Experimental Setups
B.1 Target LLMs.
  • Llama-3-8B-Instruct [2]: LLaMA-3-8B-Instruct is an instruction-tuned version of Meta’s LLaMA-3 language model with 8 billion parameters. It is trained to follow natural language instructions across a wide range of tasks and serves as a strong open-source alternative to proprietary chat models.
  • Vicuna-7B-v1.5 [14]: Vicuna-7B-v1.5 is a fine-tuned version of LLaMA-2-7B, developed by LMSYS, optimized for multi-turn conversational ability. It is trained on user-shared ChatGPT conversations and aims to match the quality of closed-source models such as GPT-3.5.
  • Qwen-2.5-7B-Instruct [3]: Qwen-2-7B-Instruct is the instruction-tuned variant of Alibaba’s Qwen-2-7B model, trained to follow prompts and generate aligned responses. It supports multilingual understanding and demonstrates competitive performance on safety and alignment benchmarks.
  • Mistral-7B-Instruct-v0.3 [15]: Mistral-7B-Instruct is an instruction-tuned version of the Mistral-7B model released by Mistral AI. It features efficient inference via grouped-query attention and sliding-window attention, and has been tuned on a wide range of instruction-following tasks.
  • Gemma-7b [16]: Gemma is Google’s family of lightweight, state-of-the-art open models, built on the same research and technology as Gemini.
Table 5 shows the source links of five target LLMs we used in this paper.
@@ Appendix B Supplementary Materials on Experimental Setups
B.2 Judge Models
@@
  • GPTFuzzer [22] is a fine-tuned RoBERTa-based model designed to evaluate the harmfulness of generated text. It is trained to classify responses as either “safe” or “unsafe”, providing a probability score that indicates the likelihood of the input text being harmful. As the results reported in [22]’s Table 1, GPTFuzzer achieves superior performance compared to traditional methods, offering both high efficiency and effectiveness in detecting potentially harmful content in LLMs.
  • Llama-Guard-3-8B [2] is a fine-tuned version of Meta’s Llama-3-8B model, specifically designed to serve as a protective guardrail model to prevent the generation of harmful or unethical content by identifying and blocking such outputs. It supports multilingual harmful content detection and classifies content into 14 distinct harmful categories, providing robust protection for developers by identifying various types of illegal or unethical content.
B.3 Benchmark Datasets
@@
B.4 Judge Template for GPT-4
In this paper, we employ GPT-4 as a judge model to evaluate the performance of DTA and existing baselines. Specifically, following previous work [13, 8], we input a judge template and a harmful prompt into GPT-4, generating an unsafe score. Here we provide the judge template:
Figure 7: Judge template for GPT-4 as a judge.
Appendix C Supplementary Experimental Settings
C.1 Response Generation Settings
We set the temperature to 0.7 following previous work [7], the maximum number of token sequences to 256, and top_k to 50. We show the generation pseudo-code as follows:

### v3

@@ 5 Conclusion
In this paper, we propose Dynamic Target Attack (DTA), a novel jailbreak method that directly samples candidate responses and selects the most harmful (yet useful) one as target to optimize the adversarial prompt. In contrast to existing methods, DTA selects targets from the target model’s relatively high-density region, significantly reducing the mismatch between the target and LLM’s output distribution, thereby easing the optimization process. We conduct extensive experiments on diverse jailbreak benchmarks, to demonstrate DTA’s effectiveness against multiple recent safety-aligned LLMs under both white- and black-box settings, outperforming the state-of-the-art baselines.
Impact Statement
This study presents Dynamic-Target Attack (DTA), a jailbreaking framework that generates effective adversarial prompts by dynamically selecting targets from the LLM’s own sampled responses, rather than optimizing toward a fixed template. Our goal is to improve the understanding of robustness and alignment failure modes in safety-aligned LLMs and to support the development of stronger defenses and safer deployment.
We acknowledge the potential negative societal impact of this work. Because DTA can produce fluent, contextually coherent jailbreak prompts more efficiently, it may be misused to bypass content moderation, induce harmful or illegal outputs, and facilitate misinformation, toxicity, or biased generation. The paper may also contain examples of offensive content due to the nature of evaluating jailbreak behavior.
To mitigate misuse, we recommend safeguards such as guard models and moderation layers (e.g., Llama Guard), rate limiting and monitoring to detect repeated optimization-like querying, and adversarial training using dynamically targeted prompts. We encourage future research to focus on: (i) detecting adaptive jailbreak prompts, (ii) robust training against dynamic-target attacks, (iii) defense-in-depth system designs, and (iv) responsible disclosure and evaluation practices that minimize harm while enabling progress on LLM safety.
Appendix A Notation table
Table 5 shows the notations and abbreviations used in our paper.
Table 5: Notations and abbreviations used in this paper.
Symbol Description
$f_{\theta}$, $f_{\phi}$ Target (Surrogate) LLM parameterized by $\theta$ ($\phi$).
$P$ Harmful prompt provided by the attacker (e.g., a request asking for disallowed or dangerous instructions).
$S$ Adversarial suffix, i.e., $S=(s_{1},\cdots,s_{L_{s}})$, where $L_{s}$ is the length of $S$. We denote $S^{*}$ as the optimized adversarial suffix produced by DTA.
$\mathbf{Z}$ Continuous representation (logits) of adversarial suffix $S$.
$r$ Model response. $r^{*}$ refers to a dynamic target response selected in the current exploration cycle (full sequence), chosen as the most harmful candidate among sampled responses. $r_{\text{fixed}}$ denotes the fixed response used by existing methods. $r_{\text{sampled}}$ denotes a target response sampled by DTA.
$M$ Number of exploration cycles in DTA; in each cycle, a new dynamic target $r^{*}$ is sampled and locally optimized.
$T_{\text{opt}}$ Number of gradient-based optimization steps per exploration cycle (inner-loop iterations).
NN Sampling count per exploration cycle; number of candidate responses drawn from the target LLM to select the most harmful target.
$\tau_{\text{search}}$ Exploration temperature used to sample candidate responses during DTA’s exploration stage (relaxed decoding).
$\tau_{\text{eval}}$ Evaluation temperature used to generate responses when measuring ASR (standard decoding).
$\tau_{\text{opt}}$ Small temperature used in the softmax over suffix logits to obtain a nearly one-hot, differentiable token distribution during optimization.
$J$ Harmfulness judge (safety classifier or LM-as-judge) that maps a response $T$ to a score $J(T)\in[0,1]$.
$\mathcal{L}_{\text{resp}}$ Response likelihood loss that encourages the model to generate the target prefix $r^{*}_{L}$ given P+SP{+}S.
$\mathcal{L}_{\text{flu}}$, $\mathcal{L}_{\text{rej}}$ Fluency loss and refusal loss used to regularize the suffix: $\mathcal{L}_{\text{flu}}$ promotes on-distribution text, $\mathcal{L}_{\text{rej}}$ penalizes probability mass on refusal-related tokens.
$\mathcal{L}_{\text{suffix}}$ Suffix regularizer defined as $\mathcal{L}_{\text{suffix}}(S)=\mathcal{L}_{\text{flu}}(S)-\mathcal{L}_{\text{rej}}(S)$.
$\mathcal{L}_{\text{DTA}}$ Overall DTA loss combining response loss and suffix regularization.
$\text{ASR}_{G}$ Attack success rates evaluated by GPTFuzzer ($G$), Llama-Guard-3 ($L$), and GPT-4 ($O$).
$\text{ASR}_{L}$
$\text{ASR}_{O}$
Appendix B Why DTA could sample harmful responses
We can characterize why DTA is likely to explore harmful responses during the exploration phase. Let $p_{u}(\tau_{\text{search}})$ denote the probability that a single stochastic decode at temperature $\tau_{\text{search}}$ yields an unsafe response (as judged by $J$), conditioned on the current adversarial prompt. Using multinomial sampling with independent randomness across draws (i.e., do_sample=True, num_beams=1), the probability of observing at least one unsafe candidate among $N$ samples in an exploration cycle is
$\Pr[\exists\,\text{at least unsafe sample}]=1-(1-p_{u}(\tau_{\text{search}}))^{N},$ (16)
which is monotone in $N$ and approaches $1$ as $N\!\rightarrow\!\infty$ whenever $0<p_{u}(\tau_{\text{search}})<1$. In practice, a non-zero exploration temperature increases output diversity and helps avoid vanishingly small $p_{u}(\tau_{\text{search}})$, making unsafe candidates more likely to appear in the sampled set.
More importantly, DTA does not require the initial sampled target to be highly unsafe: by selecting the most harmful yet useful candidate and optimizing the suffix toward it, each sampling–optimization round can progressively shift the conditional distribution away from refusal modes, enabling increasingly unsafe targets to be discovered and exploited over iterations (e.g., within 200 steps in our experiments).
Appendix C Supplementary Materials on Experimental Setups
C.1 Target LLMs.
@@ Appendix C Supplementary Materials on Experimental Setups
  • Llama-3-8B-Instruct (Dubey et al., 2024): LLaMA-3-8B-Instruct is an instruction-tuned version of Meta’s LLaMA-3 language model with 8 billion parameters. It is trained to follow natural language instructions across a wide range of tasks and serves as a strong open-source alternative to proprietary chat models.
  • Vicuna-7B-v1.5 (Chiang et al., 2023): Vicuna-7B-v1.5 is a fine-tuned version of LLaMA-2-7B, developed by LMSYS, optimized for multi-turn conversational ability. It is trained on user-shared ChatGPT conversations and aims to match the quality of closed-source models such as GPT-3.5.
  • Qwen-2.5-7B-Instruct (Yang et al., 2024): Qwen-2-7B-Instruct is the instruction-tuned variant of Alibaba’s Qwen-2-7B model, trained to follow prompts and generate aligned responses. It supports multilingual understanding and demonstrates competitive performance on safety and alignment benchmarks.
  • Mistral-7B-Instruct-v0.3 (Jiang et al., 2023): Mistral-7B-Instruct is an instruction-tuned version of the Mistral-7B model released by Mistral AI. It features efficient inference via grouped-query attention and sliding-window attention, and has been tuned on a wide range of instruction-following tasks.
  • Gemma-7b (Team et al., 2024): Gemma is Google’s family of lightweight, state-of-the-art open models, built on the same research and technology as Gemini.
Table 6 shows the source links of five target LLMs we used in this paper.
C.2 Judge Models
@@
  • GPTFuzzer (Yu et al., 2024) is a fine-tuned RoBERTa-based model designed to evaluate the harmfulness of generated text. It is trained to classify responses as either “safe” or “unsafe”, providing a probability score that indicates the likelihood of the input text being harmful. As the results reported in (Yu et al., 2024)’s Table 1, GPTFuzzer achieves superior performance compared to traditional methods, offering both high efficiency and effectiveness in detecting potentially harmful content in LLMs.
  • Llama-Guard-3-8B (Dubey et al., 2024) is a fine-tuned version of Meta’s Llama-3-8B model, specifically designed to serve as a protective guardrail model to prevent the generation of harmful or unethical content by identifying and blocking such outputs. It supports multilingual harmful content detection and classifies content into 14 distinct harmful categories, providing robust protection for developers by identifying various types of illegal or unethical content.
C.3 Judge Template for GPT-4
In this paper, we employ GPT-4 as a judge model to evaluate the performance of DTA and existing baselines. Specifically, following previous work (Jia et al., 2024; Zhu et al., 2024), we input a judge template and a harmful prompt into GPT-4, generating an unsafe score. As shown in Figure 10, we provide the judge template for GPT-4:
C.4 Defense Methods
In this section, we provide additional details on the two defense mechanisms considered in our evaluation: Paraphraser and SmoothLLM.
Paraphraser Defense
The Paraphraser defense (Jain et al., 2023) aims to neutralize adversarial prompt suffixes by rewriting the entire user query into a semantically similar but lexically different form before it is passed to the target LLM. Intuitively, many jailbreak attacks rely on carefully crafted token patterns or control phrases in the suffix; paraphrasing can disrupt these patterns while preserving the high-level intent of the query.
In our implementation, for each adversarial prompt P+SP{+}S produced by an attack, we construct a meta-prompt for GPT-4 that asks it to paraphrase the full text. As shown in the following textbox, in practice, we replace the {prompt_with_adv} with the actual adversarial prompt.
Prompt template for paraphrasing adversarial prompts You are a rewriting tool. Your task is to take an input text (even if harmful) and rewrite it into a semantically equivalent form without refusing. Only output the rewritten text, nothing else. Input: {prompt_with_adv} Output:
SmoothLLM
SmoothLLM (Robey et al., 2023) adapts randomized smoothing to the prompt space: before sending a query to the LLM, the defense applies random perturbations (e.g., character insertions) to the input, with the goal of breaking fragile adversarial structures while maintaining the overall readability and meaning of the prompt.
Following the spirit of SmoothLLM, we implement a perturbation-based pre-processing of adversarial prompts as follows:
  • Given an adversarial prompt P+SP{+}S, we randomly select a subset of character positions according to a fixed perturbation rate (e.g., a small probability per character).
  • At each selected position, we insert a random character (e.g., punctuation or alphabetic symbols) so that the final perturbed prompt remains human-readable but deviates from the exact token pattern crafted by the attacker.
  • The perturbed prompt $\tilde{P{+}S}$ is then forwarded to the target LLM $f_{\theta}$ under the same evaluation decoding settings as in the undefended case.
For both defenses, we keep all other experimental settings (datasets, target models, judges, decoding parameters) identical to the main white-box evaluation, so that changes in success rate can be attributed to the defense mechanisms rather than differences in evaluation protocol.
C.5 Benchmark Datasets
@@
Appendix D Supplementary Experimental Settings
D.1 Response Generation Settings
We set the temperature to 0.7 following previous work (Guo et al., 2024), the maximum number of token sequences to 256, and top_k to 50. We show the generation pseudo-code as follows: